OpenStudy (anonymous):

what are the factors of the polynomial 2x^2+x-10

OpenStudy (whpalmer4):

To factor the polynomial \(2x^2+x-10\), take a look at the leading coefficient, which is 2. Our first components of our factors must have a product of 2, so we'll have something like \[(2x + ____)(x + ____)\] Next, look at the final term, -10. This will be produced by the product of the two blanks. Our choices here are the factors of -10: -1,10; -2, 5; -5,2; -10,1, 1,-10;2,-5; 5,-2; 10,-1. Which choice? The third constraint is that the 2nd member of the pair * 2, plus the first member of the pair must equal 1 because that is the value of the coefficient of the middle term. 2(-2)+5 = -4+5 = 1, so that's the set to choose, giving us \[2x^2+x-10 = (2x+5)(x-2)\]If we multiply out our factorization to check (always a good idea): \[(2x+5)(x-2) = 2x^2-4x+5x-10 = 2x^2+x-10\checkmark\]